# D-Judge: How Far Are We? Accessing the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance

### v2

@@ 3. Aesthetic Appeal Discrepancy:
Aesthetic Appeal is to estimate the visual appeal and artistic quality of images, reflecting the visual attractiveness and artistic quality of images. We utilize the classical aesthetic measures NIMA (Esfandarani & Milanfar, 2018), LAION-AES (Schuhmann et al., 2022a), and QAlign (Wu et al., 2024b) as the quantitative metrics and compare the difference rates of aesthetic scores $DR_{aes}$ for the paired AI-generated and natural images.
@@ 4. Downstream Applicability Discrepancy:
This aspect is devised to investigate the practical utility of AI-generated images in downstream tasks and evaluate whether AI-generated images can have different practical utilities with realistic natural images in downstream application tasks. We mainly focus on three common downstream tasks, e.g., image recognition, object segmentation and Multimodal Visual Question Answering (VQA). For the image recognition task, we evaluate the Difference recognition rates ($DR_{re}$) between AI-generated images and natural images when using a pre-trained image recognizer like ResNet-152 (He et al., 2016) model. For the object segmentation, we use the Intersection over Union (IoU) (Everingham et al., 2010) as quantitative measures and investigate the segmentation discrepancy for AI-generated images with U2Net (Qin et al., 2020). For VQA, we design three types of questions: ‘object’ (identifying objects in the image), ‘alignment’ (determining if the image aligns with a textual description, returning ‘yes’ or ‘no’), and ‘similarity probability’ (calculating the alignment probability between text and image). Using BLIP2 (Li et al., 2023b), we evaluate the Difference Rate ($DR$) between AI-generated images and reference natural images for each question type.
@@ 5. Human Validation:
We involve human assessments to coordinate the above evaluation aspects. We develop the human assessment interface to demonstrate AI-generated images with the referenced natural images and text descriptions and collect human ratings (on a scale from 1 to 5) alongside the following three aspects: naive image quality, semantic alignment, and Aesthetic Appeal as human assessments. Human participants are provided with example images before the evaluation to understand high and low scores without being given specific numerical values, ensuring unbiased and informed ratings. Detailed information about human assessment procedures is provided in Appendix Sec. D.3.
@@
Figure 4: We evaluate nine image-generating models from 6 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
4 Evaluation
@@
  • RQ1.a: What discrepancies between I2I-guided AIGIs and their natural counterparts can be interpreted via the pixel-level quality measures?
@@
4.1 Experimental Setting
We use the described quantitative evaluation metrics, which can be divided into Full-reference and No-reference, to present the $DR$ for involved assessment dimensions (more details of all metrics are presented in Appendix Sec. C.
@@ 4.1 Experimental Setting
For full-reference metrics, which are used for pixel-level image quality, visual content distribution, and downstream object segmentation task, i.e., IoU, we directly report the calculated value to reflect the difference cause it already can present the discrepancy.
@@
$\text{{DR}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\lvert((M(\mathbf{X}_{i})-m_{\text{min}})-(M(\mathbf{N}_{i})-m_{\text{min}}))\rvert}{m_{\text{max}}-m_{\text{min}}}$ (1)
@@
4.2 RQ1: Naive Quality Results
To evaluate the image quality of AI-generated images versus natural images, we conducted a comprehensive analysis using a variety of metrics that assess the discrepancies between AIGIs and natural images across pixel-level similarity, frame-level quality and visual content distribution. Our findings reveal substantial discrepancies between AIGIs and natural images across these dimensions.
4.2.1 RQ1.a Pixel-level
@@ 4.2 RQ1: Naive Quality Results
Table 2: Full-reference Evaluation Results (I2I).
SD14 SD15 VD SD21 SDxl D·E 2
@@
4.2.2 RQ1.b Frame-level
We assessed the frame-level discrepancies of AI-generated and natural images using a suite of metrics, including DBCNN, HyperIQA, LIQE, PIQE, NRQM and QAlign. Figure 4 uses radar figures to visualize the mean difference rates from the six metrics, and indicates that there are significant quality differences between AI-generated and natural images at the frame level. These deviations are typically around 10%, irrespective of the guidance prompt used, with some instances showing quality differences as high as 40%. This underscores the considerable gaps between AIGIs and natural images in terms of global image quality. Moreover, the results of NRQM reveal that AI-generated images significantly differ from natural images in terms of naturalness. Specifically, AI-generated images show notable deviations from natural images in visual realism and adherence to natural scene statistics. The NRQM difference rates range from 5% to 32%, with images generated by D·E 3 exhibiting particularly high rates, underscoring the challenges AI-generated images encounter in replicating the inherent naturalness of real-world scenes. More detailed results with more metrics are presented in Appendix Sec. D.1 Figure 12.
Figure 5: The FID (left, ↓) and Inception Score (right, ↑) of AI-generated images from different source models and guidance.
4.2.3 RQ1.c Visual Content Distribution
@@
In conclusion, our evaluation of image quality across pixel-level, frame-level, and content distribution-level reveals significant disparities between AI-generated and natural images. These differences highlight the current limitations of AIGC technologies in matching the visual quality and realism of natural images.
4.3 RQ2: Alignment Results
We calculate the difference rate of CLIP Scores ($DR_{CS}$) to assess how well the AIGIs align with natural ones in terms of semantic alignment across different guidance types: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I). And we find that AIGIs often struggle to maintain semantic alignment as natural ones, especially the image-only reference provided. The results, shown in Table 3, indicate that the $DR_{CS}$ range from 4.02% to 36.69%. We also find that even if the AIGIs can get a higher CS sometimes, especially with text and image-guided, they are looking stranger to humans. This suggests that current AIGC models often struggle to maintain high fidelity to the provided references, leading to variations that may impact the intended alignment. Our findings underscore the challenges that these models face in consistently generating content to accurately reflect the specified input conditions, highlighting both their strengths and limitations in real-world applications.
Table 3: The results of $DR_{CS}$ (%).
@@ 4.4 RQ3: Aesthetic Appeal
We assessed the aesthetic appeal of AI-generated and natural images using NIMA and LAION-AES metrics and found that AI-generated images generally fall short of natural images in terms of aesthetic quality. Figure 6 presents the aesthetic difference rate ($DR_{aes}$) results, which reveal that AI-generated images have noteworthy discrepancies compared to natural ones, with $DR_{aes}$ ranging from 3.51% to 27.66%. These findings indicate that while AI models can produce visually appealing images, AIGIs’ often lack the nuanced artistic quality and emotional impact that are characteristic of natural images.
Figure 6: The $DR_{aes}$ between AIGIs and natural ones.
4.5 RQ4: Applicability Results
When evaluating the applicability of AI-generated images, we focused on three crucial downstream tasks: image recognition, semantic segmentation and VQA. Our findings indicate significant discrepancies between the performance of AIGC images and natural images in these tasks.
Image Recognition:
We measured the classification Difference Rate (DR) (%) to identify the differences in predicted labels between AIGC images and their natural counterparts. The results in Table 8 showed a substantial gap, with the DR ranging from 27.46% to 94.29%, indicating that the generated images can’t accurately present the object as expected.
Semantic Segmentation:
We evaluated the performance using the Intersection over Union (IoU) metric and show the discrepancy results in Table 9. The $IoU$ values (higher is better) for AIGC images varied between 0.33 and 0.82, indicating a wide range of segmentation between AIGC images and their natural counterparts.
Visual Question Answering:
Further, we apply these AI-generated image to the VQA task by BLIP-2 (Li et al., 2023b) involved three different types of questions: Object, Alignment (Align), and similarity probability (Prob.). Table 4 summarizes the discrepancies (%) between natural and AI-generated images across various questions. For Object, the highest discrepancies occurred in the I2I setting, with SD14 and SD15 reaching 96.41% and 96.95%, respectively, while T2I discrepancies ranged from 64.07% (DALL·E 3) to 88.43% (DFGAN). Align questions showed lower discrepancies, with T2I values ranging from 8.55% (DFGAN) to 21.18% (DALL·E 3) and I2I values from 7.87% (VD) to 10.07% (SD15). For similarity probability (Prob.), T2I discrepancies varied between 37.31% (DALL·E 2) and 62.69% (DFGAN), while I2I results ranged from 32.81% (DALL·E) to 61.56% (SD15).
Overall, our findings highlight significant limitations of AI-generated images in tasks such as image recognition, semantic segmentation, and visual question answering (VQA). While AI-generated images perform comparably to natural images in simpler tasks, such as binary text-image alignment (e.g., responding with “yes” or “no”), they fall short in more complex scenarios, including accurately estimating text-image alignment probabilities, identifying specific objects, and segmenting object areas. These shortcomings limit their broader applicability in diverse downstream tasks compared to natural images. Our results underscore the need for continued research and development to improve the reliability and accuracy of AIGC models, enabling their effective deployment in real-world applications.
Table 4: The Difference Rate (DR) of VQA tasks.
@@ 4.5 RQ4: Applicability Results
4.6 Human Validation
To rigorously assess the differences between AI-Natural images and to evaluate the effectiveness of existing metrics in measuring AIGC image quality, we conducted human evaluations focusing on three key aspects: Image Quality, Alignment, and Aesthetic Appeal.
  • Image Quality: Participants assessed the overall quality of the images, considering factors such as clarity, detail, and the presence of artifacts.
  • Alignment: This aspect measured how well the generated images matched the given prompts.
  • Aesthetic: The aesthetic appeal of the images was rated based on visual attractiveness, composition, and overall artistic quality.
We excluded results for Applicability since this dimension inherently involves human-generated ground truth labels. Participants rated images on a scale from 1 (poor) to 5 (excellent). Our human evaluation results are shown in Table 5 and Figure 7 reveal significant discrepancies between AI-generated images and natural images across all three assessed aspects: quality, alignment, and aesthetic appeal. AIGC images consistently received lower scores across all three dimensions; AI-generated images generally received lower scores, often between 2 and 3, compared to natural images. This emphasizes the need for further advancements to narrow these gaps. These findings underscore the importance of continued research and development in AI image generation to bring AIGC images closer to the standards of natural images.
Moreover, the result in Figure 7 from human validation indicates that traditional quantitative metrics often fail to align with human preferences, suggesting that they are not fully compatible with assessing AIGC images. This highlights the necessity for developing new measures tailored specifically to the unique characteristics of AI-generated content.
Table 5: Human score results across three aspects.
SD21 SDxl VD D·E 2
@@ 4.6 Human Validation
Figure 7: The difference rate provided by human & quantitative statistics, where ‘H’ represents human and ‘Q’ represents quantitative.
4.7 Discrepancies Across Categories
@@
To analyze the differences between AI-generated and natural images across various categories, we evaluated the discrepancy of AIGI from eight categories (‘person’, ‘animal’, ‘indoor’, ‘outdoor’, ‘vehicle’, ‘food’, ‘sports’, ‘accessory’) across three key aspects: ➊ Naive Quality, including pixel-level similarity (SSIM), frame-level quality (CLIPIQA), and content distribution (FID, CLIP-FID, and Inception Score); ➋ Alignment, measured by CLIP Score; and ➌ Aesthetic Quality, assessed by LAION-AES. The results in Figure 8 reveal substantial variability across different categories, even for the AIGI from the same model (please refer to Appendix Sec. D.2, Figures 16 and 17 for integrated details). This could be attributed to the significant imbalance in the distribution of training data across these categories (Schuhmann et al., 2022b; Aghajanyan et al., 2023).
@@
Figure 8: The difference rate on food and sport subcategories across quality, alignment, and aesthetic.
5 Conclusion and Discussion
In this paper, we develop a systematic and comprehensive D-Judge to evaluate the discrepancies between AI-Natural images using the constructed D-ANI dataset. Our findings reveal significant disparities in naive image quality, semantic alignment, aesthetics, and downstream task applicability, especially in the low-level and content distribution-level image quality and fine-grained downstream tasks, like object recognition and object VAQ. While AIGC models have achieved impressive innovations, AIGIs still lag behind natural images in these critical areas. Besides, human evaluations also reveal lower scores for AIGC in terms of visual quality, semantic alignment, and aesthetics, emphasizing the gaps between AIGIs and natural images. Finally, the poor utility of AIGIs in downstream tasks indicates practical application challenges for AIGC.
This study also has limitations, including reliance on specific quantitative metrics and models available at the time. Future studies should incorporate more generative models, larger datasets, and sophisticated metrics to capture further discrepancies. Addressing these limitations and improving models will help reduce the perceptual gaps between AIGC images and natural images, facilitating better integration into real-world applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@ 3. Aesthetic Appeal Discrepancy:
Aesthetic Appeal is to estimate the visual appeal and artistic quality of images, reflecting the visual attractiveness and artistic quality of images. We utilize the classical aesthetic measures NIMA (Esfandarani and Milanfar, 2018), LAION-AES (Schuhmann et al., 2022), and QAlign (Wu et al., 2024) as the quantitative metrics and compare the difference rates of aesthetic scores $DR_{aes}$ for the paired AI-generated and natural images.
@@ 4. Downstream Applicability Discrepancy:
This aspect is devised to investigate the practical utility of AI-generated images in downstream tasks and evaluate whether AI-generated images can have different practical utilities with realistic natural images in downstream application tasks. We mainly focus on three common downstream tasks, e.g., image recognition, object segmentation and Multimodal Visual Question Answering (VQA). For the image recognition task, we evaluate the Difference recognition rates ($DR_{re}$) between AI-generated images and natural images when using a pre-trained image recognizer like ResNet-152 (He et al., 2016) model. For the object segmentation, we use the Intersection over Union (IoU) (Everingham et al., 2010) as quantitative measures and investigate the segmentation discrepancy for AI-generated images with U2Net (Qin et al., 2020). For VQA, we design three types of questions: ‘object’ (identifying objects in the image), ‘alignment’ (determining if the image aligns with a textual description, returning ‘yes’ or ‘no’), and ‘similarity probability’ (calculating the alignment probability between text and image). Using BLIP2 (Li et al., 2023), we evaluate the Difference Rate ($DR$) between AI-generated images and reference natural images for each question type.
@@ 5. Human Validation:
We involve human assessments to coordinate the above evaluation aspects. We develop the human assessment interface to demonstrate AI-generated images with the referenced natural images and text descriptions and collect human ratings (on a scale from 1 to 5) alongside the following three aspects: naive image quality, semantic alignment, and Aesthetic Appeal as human assessments. Human participants are provided with example images before the evaluation to understand high and low scores without being given specific numerical values, ensuring unbiased and informed ratings. Detailed information about human assessment procedures is provided in Sec. 4.6.
@@
Figure 4. We evaluate nine image-generating models from 6 aspects. The numerical values in the radar chart represent the mean difference rate of each model.
4. Evaluation
@@
  • RQ1.a: What discrepancies between I2I-guided AIGIs and their natural counterparts can be interpreted via the structure-level quality measures?
@@
4.1. Experimental Setting
We use the described quantitative evaluation metrics, which can be divided into Full-reference and No-reference, to present the $DR$ for involved assessment dimensions (more details of all metrics are presented in Appendix. C.
@@ 4.1. Experimental Setting
For full-reference metrics, which are used for structure-level image quality, visual content distribution, and downstream object segmentation task, i.e., IoU, we directly report the calculated value to reflect the difference cause it already can present the discrepancy.
@@
(1) $\text{{DR}}=\frac{1}{n}\sum_{i=1}^{n}\frac{\lvert((M(\mathbf{X}_{i})-m_{\text{min}})-(M(\mathbf{N}_{i})-m_{\text{min}}))\rvert}{m_{\text{max}}-m_{\text{min}}},$
@@
4.2. RQ1: Naive Quality Results
To evaluate the image quality of AI-generated images versus natural images, we conducted a comprehensive analysis using a variety of metrics that assess the discrepancies between AIGIs and natural images across structure-level similarity, frame-level quality and visual content distribution. Our findings reveal substantial discrepancies between AIGIs and natural images across these dimensions.
4.2.1. RQ1.a Structure-level
@@ 4.2. RQ1: Naive Quality Results
Table 3. Full-reference Evaluation Results (I2I).
Metrics SD14 SD15 VD SD21 SDxl D·E 2
@@
4.2.2. RQ1.b Frame-level
We assessed the frame-level discrepancies of AI-generated and natural images using a suite of metrics, including DBCNN, HyperIQA, LIQE, PIQE, NRQM and QAlign. Figure 4 uses radar figures to visualize the mean difference rates from the six metrics, and indicates that there are significant quality differences between AI-generated and natural images at the frame level. These deviations are typically around 10%, irrespective of the guidance prompt used, with some instances showing quality differences as high as 40%. This underscores the considerable gaps between AIGIs and natural images in terms of global image quality. Moreover, the results of NRQM reveal that AI-generated images significantly differ from natural images in terms of naturalness. Specifically, AI-generated images show notable deviations from natural images in visual realism and adherence to natural scene statistics. The NRQM difference rates range from 5% to 32%, with images generated by D·E 3 exhibiting particularly high rates, underscoring the challenges AI-generated images encounter in replicating the inherent naturalness of real-world scenes. More detailed results with more metrics are presented in Appendix. D.1 Figure 13.
Figure 5. The FID (left, ↓) and Inception Score (right, ↑) of AI-generated images from different source models and guidance.
4.2.3. RQ1.c Visual Content Distribution
@@
In conclusion, our evaluation of image quality across structure-level, frame-level, and content distribution-level reveals significant disparities between AI-generated and natural images. These differences highlight the current limitations of AIGC technologies in matching the visual quality and realism of natural images.
4.3. RQ2: Alignment Results
We calculate the difference rate of CLIP Scores ($DR_{CS}$) to assess how well AI-generated images (AIGIs) align with natural images in terms of semantic consistency across three types of guidance: Text-to-Image (T2I), Image-to-Image (I2I), and Text-and-Image-to-Image (TI2I). Our results show that AIGIs often struggle to achieve the same level of semantic alignment as natural images, particularly when guided solely by image input. As presented in Table 4, the $DR_{CS}$ ranges from 4.02% to 36.69%. This suggests that current AIGIs often face difficulties in preserving high-fidelity alignment with the provided inputs, leading to outputs that deviate from natural counterparts. These findings highlight the ongoing challenges that AIGIs encounter in faithfully and consistently capturing the intended semantics, underscoring both their capabilities and limitations in real-world scenarios.
Table 4. The results of $DR_{CS}$ (%).
@@ 4.4. RQ3: Aesthetic Appeal
We assessed the aesthetic appeal of AIGIs and natural images using two widely adopted metrics: NIMA and LAION-AES. Our results indicate that AI-generated images generally fall short of natural images in terms of aesthetic quality. As shown in Figure 6, the aesthetic difference rate ($DR_{aes}$) ranges from 3.51% to 27.66%, highlighting noticeable perceptual discrepancies between AIGIs and their natural counterparts. These findings suggest that although current AI models are capable of generating visually appealing content, AIGIs often lack the nuanced artistic detail and emotional resonance that are characteristic of natural images. This aesthetic gap underscores the limitations of existing generative systems in replicating the subtle visual cues and expressive depth found in human-created imagery.
Figure 6. The $DR_{aes}$ between AIGIs and natural ones.
Table 5. Image Recognition Discrepancy Rate ($DR_{re}$).
GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 68.60 85.46 66.14 65.50 68.12 64.88 64.88 65.36 69.26
I2I - - 94.29 94.07 40.90 83.30 27.96 47.98 -
TI2I - - 61.62 59.71 60.05 55.54 27.46 - -
Table 6. IoU of Object Segmentation.
GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 0.23 0.24 0.26 0.25 0.26 0.25 0.26 0.26 0.28
I2I - - 0.38 0.38 0.69 0.50 0.82 0.42 -
TI2I - - 0.43 0.43 0.29 0.53 0.82 - -
4.5. RQ4: Applicability Results
We evaluated the applicability of AIGIs across three representative downstream tasks: image recognition, semantic segmentation, and visual question answering (VQA). Our results reveal substantial performance discrepancies between AI-generated and natural images across all tasks, raising concerns about the practical utility of current AIGIs in real-world applications.
4.5.1. Image Recognition
We quantified recognition inconsistencies using the Difference Rate ($DR_{re}$, %) of predicted labels. As shown in Table 5, $DR_{re}$ values ranged from 27.46% to 94.29%, indicating that AI-generated images often fail to accurately represent the intended visual concepts, leading to incorrect or unstable recognition outcomes.
4.5.2. Semantic Segmentation
Semantic segmentation performance was evaluated using Intersection over Union (IoU), computed based on segmentation results produced by U2Net (Qin et al., 2020). Table 6 reports IoU values ranging from 0.33 to 0.82, with a mean IoU of 0.40. These results suggest that generated images frequently suffer from imprecise object boundaries and incorrect spatial configurations, limiting their usability for structure-level understanding tasks.
4.5.3. Visual Question Answering
We further assessed AI-generated images in the context of visual question answering using BLIP-2 (Li et al., 2023), focusing on three types of questions: Object Recognition, Semantic Alignment (Align), and Similarity Probability (Prob.) (The prompt template used here please refer to Appendix D.2). As summarized in Table 7, object-based questions exhibited the highest discrepancy rates, particularly under the I2I setting, reaching up to 96.41% (SD14) and 96.95% (SD15). Binary alignment questions (only answer ”yes” or ”no”) yielded lower discrepancies, with an average $DR$ of 11.55%, while similarity-based probabilistic questions underscoring substantial discrepancies compared to natural images, with an average $DR$ of 50.41%.
Summary.
Although AI-generated images perform relatively well as natural ones on simpler binary tasks, their effectiveness significantly deteriorates in more complex settings that require fine-grained visual understanding, such as object recognition, spatial reasoning, and nuanced semantic interpretation. These limitations currently hinder the broader adoption of generative images in downstream vision tasks. Our findings highlight an urgent need for further research into improving the structural fidelity and semantic coherence of generated content to enable trustworthy and task-reliable deployment.
Table 7. The Difference Rate (DR,%) of VQA tasks.
@@ 4.5.3. Visual Question Answering
4.6. Human Validation
To rigorously examine the gap between AI-generated images and human expectations, as well as to evaluate the reliability of existing metrics in assessing AI-generated images, we conducted a comprehensive human evaluation study. Participants were asked to score images across three key dimensions: Image Quality, Alignment, and Aesthetic Appeal.
  • Image Quality: Evaluates the overall visual fidelity of the image, including clarity, sharpness, and the presence of visual artifacts.
  • Alignment: Assesses the semantic consistency between the generated image and its corresponding prompt, which could be textual, visual, or a combination of both.
  • Aesthetic Appeal: Measures the artistic and visual attractiveness of the image, considering aspects such as composition, color balance, and overall presentation.
Participants were instructed to rate each image on a 5-point scale (1 = poor, 5 = excellent) for each of the three evaluation aspects using a radio-button interface.
4.6.1. Interface for Human Evaluation
The evaluation was implemented through a custom-designed interface (please refer to Appendix. D.3, Figure 15). This interface enabled participants to view AI-generated images alongside their reference prompts and assign scores accordingly. The prompts used in the evaluation were drawn from the D-ANI dataset and were provided in three modalities: (1) text-only, (2) image-only, and (3) text–image combinations. The evaluated images were generated by four state-of-the-art models: SD21, SDxl, Versatile Diffusion, and DALL·E 2.
To support score calibration and mitigate evaluation bias, we incorporated qualitative reference examples into the interface (please refer to Appendix. D.3, Figure 16). These examples illustrated typical characteristics of both strong and weak performance across the three evaluation aspects, without disclosing specific numeric scores, thereby reducing potential anchoring effects.
4.6.2. Evaluation Process
A total of 58 volunteers participated in the evaluation. Each participant rated a curated set of images based on the provided prompts and scoring criteria. To assess the consistency and reliability of the responses, we included duplicated image entries within the evaluation set. By comparing the scores assigned to these repeated images, we identified inconsistent raters and excluded their data from further analysis. After this filtering process, 47 participants’ responses were retained as valid samples for final analysis.
4.6.3. Participant Demographics
The participants ranged in age from 19 to 46 years, with a gender distribution of 60% male and 40% female. The majority held at least an undergraduate degree, ensuring a well-educated evaluation cohort. In addition to scoring data, we collected basic demographic information and invited participants to provide open-ended feedback on their evaluation experience, particularly regarding the perceived differences between AI-generated and natural images.
4.6.4. Results and Observations
Quantitative results from the human study are summarized in Table 8, and visualized in Figure 7. The findings reveal a pronounced gap between AI-generated images and human expectations across all three evaluation dimensions. Specifically, AIGC images consistently received modest scores—typically between 2 and 3—indicating perceptible shortcomings in terms of visual quality, semantic fidelity, and aesthetic appeal. These results highlight the need for continued advancements in image generation technologies to better align with human perceptual standards.
Moreover, the analysis reveals that conventional quantitative metrics often fail to reflect human perceptual judgments (see Figure 7), suggesting that such metrics may be insufficient for evaluating AIGC images. This underscores the importance of developing new, human-aligned evaluation measures tailored to the unique characteristics of AI-generated visual content.
Table 8. Human scores across three aspects.
Aspect SD21 SDxl VD D·E 2
@@ 4.6.4. Results and Observations
Figure 7. The difference rate provided by human & quantitative statistics, where ‘H’ represents human and ‘Q’ represents quantitative.
4.7. Discrepancies Across Categories
@@
To systematically investigate the differences between AI-generated images (AIGI) and natural images across various content categories, we evaluate discrepancies in eight representative categories: person, animal, indoor, outdoor, vehicle, food, sports, and accessory. The evaluation is conducted along three critical dimensions: ➊ Naive Quality, assessed via structure-level similarity (SSIM), frame-level quality (CLIPIQA), and content distribution metrics (FID, CLIP-FID, and Inception Score); ➋ Alignment, measured using CLIP Score; and ➌ Aesthetic Quality, evaluated through LAION-AES.
Representative results are illustrated in Figures 8 and 9. Specifically, Figure 8 presents the difference rate ($DR$) across CLIPIQA, CLIP Score, and LAION_AES on the food and sports subcategories, reflecting variations in visual quality, semantic alignment, and aesthetic perception. Figure 9 further provides FID and Inception Score comparisons that reveal discrepancies in content distribution for the person and accessory categories.
These results demonstrate that the extent of divergence between AI-generated and natural images varies significantly across categories, even when produced by the same generative model. A potential cause of this variability is the inherent imbalance in training data distributions across different semantic categories (Schuhmann et al., 2022; Aghajanyan et al., 2023). Additional results and visualizations are provided in Appendix D.4, Figures 17 and 18, offering a comprehensive view of these discrepancies.
Notably, Figures 18 suggest that images generated by D·E 3 display the most pronounced deviations from natural references across multiple metrics. This could be attributed to D·E 3’s inherent tendency to generate visually appealing but stylized or surreal imagery, which differs from the more realism-oriented outputs of other models (see  B, Figure 10 for visual comparisons).
@@
Figure 8. The difference rate on food and sport subcategories across quality, alignment, and aesthetic.
Figure 9. The FID and Inception Score (IS) of AI-generated images from different source models and categories from content distribution level.
5. Conclusion
In this paper, we present a comprehensive D-Judge based on the constructed D-ANI dataset to evaluate the discrepancies between AI-generated and natural images. The results reveal substantial gaps across key dimensions such as image quality, semantic alignment, and downstream applicability. Despite recent progress in AIGC models, human evaluations consistently show that AI-generated images underperform in visual fidelity, semantic relevance, and aesthetic quality. Our benchmark also yields actionable guidance for improving model design, such as enhancing semantic consistency and prompt-image alignment, particularly in I2I tasks. Promising directions include reference-aware attention, perceptual alignment losses, and joint image-text encoding.
Limitation and Future Work. This study has certain limitations, including its reliance on quantitative metrics and a limited set of generative models. Future work should incorporate more diverse and balanced datasets, a broader range of contemporary models, and task-aware evaluation protocols to uncover deeper discrepancies. Beyond evaluation, D-Judge can also serve as a training signal—for instance, as a reward module in reinforcement learning or preference-based fine-tuning—to jointly optimize semantic alignment and downstream usability. We will release our dataset and code to support such use cases. Addressing these limitations and advancing model architectures will be essential for narrowing the perceptual and functional gaps between AI-generated and natural images, ultimately enabling their trustworthy deployment in real-world scenarios.
Acknowledgments
This work is supported by Guangdong Basic and Applied Basic Research Foundation (2023A1515012848) and CCF-DiDi GAIA Collaborative Research Funds. This research is supported by A*STAR, CISCO Systems (USA) Pte.Ltd and National University of Singapore under its Cisco-NUS Accelerated Digital Economy Corporate Laboratory (Award I21001E0002). We would like to express our gratitude to all the volunteers who participated in our human evaluation experiments.